Brian Canning Co., which sells canned corn, uses an operation costing system. Cans of corn are classified as either sweet or regular, depending on the type of corn used. Both types of corn go through the separating and cleaning operations, but only regular corn goes through the creaming operation. During January, two batches of corn were canned from start to finish. Batch X consisted of 1,700 pounds of sweet corn, and Batch Y consisted of 1,600 pounds of regular corn. The company had no beginning or ending Work-in-Process Inventory. The following cost information is for the month of January:

Batch Cost Batch Size
Raw sweet corn X $ 11,730 1,700 lbs
Raw regular corn Y 6,320 * 1,600 lbs
Separating department costs 4,455
Cleaning department costs 2,970
Creaming department costs 1,120


*Includes $2,100 for cream.

Required:

1. Compute the unit cost for sweet corn and regular corn.
Answer #1
Raw sweet corn = (11730/1700) = $ 6.90/.
Raw regular corn = (6320/1600) = $ 3.95/.
Total pounds for Separating and Cleaning dept = (1700+1600) = 3,300
Total pounds for creaming dept = 1,600
Separating department cost = (4455/3300) = $ 1.35
Cleaning department cost = (2970/3300) = $ 0.845
Creaming department cost = (1120/1600) = $ 0.70
Total product cost per pound :-
Sweet corn = (6.9+1.35+0.845) = $ 9.10/. Approx.
Regular corn = (3.95+1.35+0.845+0.70) = $ 6.85/. Approx.